Meenakshi: ‘How old are you ,Mom’
Mom: ‘Let me think, Meenakshi, well our three ages add up exactly seventy years’
Meenakshi: That it a lot, is not it? And how old are you Dad?
Dad: Just Six times as old as you, my daughter.
Meenu: Shall I ever be half as old as you, Dad?
Dad: ‘Yes , Meenakshi, when that happens our three ages will add up to exactly twice as much as today’

Answer by ankor@dixie-net.com(22740)

This problem will not solve with given values, I am going to change it so it does.
:
Meenakshi: ‘How old are you ,Mom’
Mom: ‘Let me think, Meenakshi, well our three ages add up exactly seventy-two years’
Meenakshi: That it a lot, is not it? And how old are you Dad?
Dad: Just Six times as old as you, my daughter.
Meenu: Shall I ever be half as old as you, Dad?
Dad: ‘Yes , Meenakshi, when that happens our three ages will add up to exactly twice as much as today’
:
Let m = Mom's present age
Let n = Meena's age
Let d = Dad's age
:
Write an equation for each statement
:
" our three ages add up exactly seventy-two years’"
m + n + d = 72
:
"Dad: Just Six times as old as you, my daughter."
d = 6n
:
"Shall I ever be half as old as you, Dad. when that happens our three ages will add up to exactly twice as much as today’
:
From this statement we know that the total of the 3 people increases by 72 yrs
Therefore each person's age has to increase by 72/3 = 24 yrs
now the equation is:
d + 24 = 2(n+24)
d + 24 = 2n + 48
d = 2n + 48 - 24
d = 2n + 24
Replace d with 6n
6n = 2n + 24
6n - 2n = 24
4n = 24
n = 24/4
n = 6 yrs old is Meena's present age
then
6(6) = 36 yrs is Dad's age
and
6 + 36 + m = 72
m = 72 - 42
m = 30 yrs is Mom's present age
:
:
Check: 36 + 30 + 6 = 72